Consider the following case. Of 200 people in a survey, 120 are men and 80 are women. Of the 120 men, 40 own stock in a US corporation, while of the 80 women, 20 are stockholders. Determine the probability of selecting one person at random who is a man and owns stock.


1
Expert's answer
2021-09-12T16:41:47-0400

"P_{Men}= \frac{120}{200}=\frac{3}{5}\\
P_{Women}= \frac{80}{200}=\frac{2}{5}\\
P_{Men \space with \space stock}= \frac{40}{120}=\frac{1}{3}\\
P_{Women \space with \space stock}= \frac{20}{80}=\frac{1}{4}\\"

Selecting one person at random who is a man and owns stock

"P=\frac{3}{5}* \frac{1}{3}= \frac{1}{5}"